Dear all, please find below a short summary of the activities of the 27th related to the AO system.
People at LBTO:G. Brusa people working from remote:A. Riccardi, M. Xompero (VNC+Skype) March 28th
Total # of person: 3
Summary:First night in about one week with virtually no problems as far as operation of the ASM! The mirror ran the entire night smoothly and without faults, there still were problems of GUIs 'dying' at times but without compromising the ASM operation. In one instance the AO Arbitrator stopped with the effect of not being able to communicate the current elevation, in this case there were no problems but the mirror could have reset during the next change in elevation if the remote operator (Armando) had not caught this event in time. In a second instance (end of night) the AO Arbitrator GUI did not allow to set the ASM in rest position, the command had to be issued from the AdSec Arbitrator GUI (lower level process).
Short log:UT 01:57 Opened dome open with telescope horizon pointing.
UT 01:58 Setting ASM while slewing towards the first object.
UT 02:20 Coma corrected ‘by hand’ by translating M2 with the hexapod by ~2mm. There is a large amount of astigmatism (visible by going slightly out of focus), the circle of ‘least confusion’ has a diameter of about 2” and the seeing (as provided by the DIMM) should be ony 1”. From this we derive a peak of astigmatism of ~ 10um wf which corresponds to 5um p surface or 2um rms. NOTE: The primary was much cooler than ambient (-4C vs. 2C).

UT 02:06 actuator dsp# 468 had a sudden jump of ~0.7um, if it had the ICC (control current integrator) the force would have jumped beyond the acceptable value (or very close). UT 02:40 pointing a different star, we found that the astigmatism is not there anymore…, the coma is still ok. El.~25 deg.
UT 02:50 Elevation 62 deg, wind 5 m/s, astigmatism is back. Cooling inlet 4.8, outlet 6.7, hub 2.1, cp 4.8, rb 4.2.
UT 02:59 the active correction put ~-150nm (rms?) of Z6, which corrected the astigmatism but appears to be small wrt to what was estimated ‘by eye’ on the previous star (UT 02:20).
UT 03:32 Troubleshooting the active optics corrections, it appears that there are some problems with the stability of the loop with off-axis guiding.
UT 03:47 some changes in the code (a sign) appear to have fixed the problem of guiding off-axis, now checking a different angles in the active optics field.
UT 04:53(?) the GCS code does not seem to work, it is decided to switch to using directly the IDL code to see if there are any differences.
UT 04:50 some parameters needed to be adjusted in the GCS code (probe rotation) and now it appears that the off-axis guiding is working. The seeing however has deteriorated.
UT 04:53 tracking object at 56 deg elevation, wind 2 m/s. Cooling flow 10 l/min. inlet 5.3, outlet 7.3, hub 2.0, cp 5.3, rb 4.4.
UT 05:37 tracking object at 65 deg elevation, wind 1m/s. Cooling flow 10 l/min. Inlet 5.7, outlet 7.7, hub 1.8, cp 5.8, rb 4.9, air temperature in the dome is reported to be 1.8.
UT 06:05 Marco connected via VNC, decided to load the default flat (20100316_185538_close_loop_200modes_a_mean.sav), for some reason at the beginning of the night a different one was loaded.
UT 08:31 starting with field aberration for active optics. ASM is working well, except actuator dsp#366 with a force at -0.6N (limit at -0.78N).
UT 09:06 sometime in the past half hour actuator dsp#366 went back to ‘normal’ current, (thanks, probably, to a jump back to its original readings).
UT 09:23 tracking object at 70 deg elevation, wind 2 m/s. Coolant flow 10 l/min. inlet 4.9C, outlet 6.8, hub 4.1, cp 4.8, rb 5.6. Now the cooling temperature (inlet) appears to the tracking the dome temperature (hub).
UT 11:57 tracking object at 55 deg elevation, wind 1 m/s. Cooland flow 10 l/min. Inlet 4.6, outlet 7.0, hub 4.8, cp 5.5, rb 5.6.
UT 12:46 first attempt to rest failed, AdOptGui reports “NOT PROCESSED”. Decided to proceed as yesterday using the AdSec Control GUI, it was successful. AdSec Control GUI reports: SHELL RIP, TSS OFF, POWER ON. The local ‘ear’ at 7 o’clock is still visible.
Ciao,Guido

-- ArmandoRiccardi - 02 Apr 2010
Topic revision: r1 - 02 Apr 2010, ArmandoRiccardi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback